# KBAda: Efficient Self Adaptation on Specific Knowledge Bases

### v1

@@ Abstract
Humans can utilize techniques to quickly acquire knowledge from specific materials in advance, such as creating self-assessment questions, enabling us to achieving related tasks more efficiently. In contrast, large language models (LLMs) usually relies on retrieval-augmented generation to exploit knowledge materials in an instant manner, or requires external signals such as human preference data and stronger LLM annotations to conduct knowledge adaptation. To unleash the self-learning potential of LLMs, we propose KBAda, an approach designed for efficient adaptation to downstream tasks involving knowledge bases. Our method utilizes iterative training with self-annotated data such as Q&A pairs and revision suggestions, enabling the model to grasp the knowledge content efficiently. Experimental results on multiple datasets demonstrate the effectiveness of our approach, significantly boosting model performance in downstream tasks that require specific knowledge at a low cost. Notably, our approach achieves over 90% of the performance improvement that can be obtained by using GPT-4-turbo annotation, while relying entirely on self-supervision. We release our experimental data, models, and process analyses to the community for further exploration 111 https://github.com/thunlp/KBAda.
KBAda: Efficient Self Adaptation on Specific Knowledge Bases
Zheni Zeng1*, Yuxuan Chen1*, Shi Yu1, Yukun Yan1🖂, Zhenghao Liu2, Shuo Wang1, Xu Han1, Zhiyuan Liu1, Maosong Sun1, 1Tsinghua University, Beijing, China 2Northeastern University, Shenyang, China yanyk.thu@gmail.com
@@ 1 Introduction
Figure 1: Comparison between different knowledge adaptation methods. KBAda requires no external signals for tuning, and improves by itself in both tuning and inference stages.
@@
Inspire by the above analogy, we purpose KBAda, a highly efficient self adaptation approach tailored to specific knowledge bases (KBs) comprising self-improvement learning. Generally, we tune the model in a completely self-supervised manner, and maximize the tuned model’s potential in inference. As shown in Fig. 1, for self annotation, we organize the original KB materials and conduct self annotation to get initial Q&A data pairs that can cover various downstream task scenarios. For iterative tuning, we require the model to check its own responses as the learning progresses, and help the model to modify the common mistakes at the current stage. Meanwhile, we conduct targeted inference in which strategies such as query expansion and confidence verification are adopted to refine the response.
Experiments on fact QA, long-form QA, and professional field test have shown the effectiveness of our method across different backbone models. With a low cost, KBAda masters the general knowledge content from the KB and achieves significant performance improvements. Side experiments including ablation studies and performance curve analyses identify the most efficient self-annotated data amount and optimal training volumes, offering valuable guidance for effectively applying KBAda in practical scenarios.
Our main contributions are as follows: (1) We propose KBAda, a novel method for autonomous LLM adaptation tailored to knowledge bases. KBAda is the first approach to help LLMs perform KB adaptation relying entirely on self annotations; (2) We provide empirical insights into efficient self adaptation to KBs, offering practical parameters and settings for deploying KBAda; (3) We conduct a comprehensive analysis of the proposed self adaptation framework. Through a range of evaluation metrics and case studies, we identify the effectiveness of KBAda and discuss the current limitations of our approach, highlighting areas for future improvement.
@@ 2 Related Work
Figure 2: Details for the KBAda framework. (a) Schematic and data composition for the whole process; (b) Instances for different annotation strategies and tasks.
@@ 4.3 Other Settings
Table 1: KBAda adaptation experiments on LooGLE, ASQA, and JEC-QA.
@@ 4.4 Result Analysis
Main Experiments. The main results for our experiments are shown in Table 1. We provide the GPT-series results, the initial version and the self-adapted version of both MiniCPM-2.4B and LLaMA-3.1-8B-Instruction on the three dataset. Overall, comparing the “Ours” lines with the corresponding original model, we can see that KBAda ensures an obvious improvement on most of the metrics, regardless of the dataset and the backbone model. This shows the generalization and effectiveness of our method.
@@
Figure 3: The impact of training amount on LooGLE performance.
@@
Figure 5: Cases for KBAda and baseline comparison. We display the translation for the Chinese JEC-QA task. The bold text and underlined text providing correct and wrong information for the QA process.
@@ 4.5 Case Studies
We display typical cases in Fig. 5 to explain the specific usefulness of our strategies. “Baseline” refers to MiniCPM-2.4B and “ours” refers to the adapted version of it. Overall, KBAda achieves a general grasp of current KB, a better knowledge answering, and a reasonable confidence verification.
@@ 5 Discussion and Limitations
In this paper, we introduced KBAda, a highly efficient self-adaptive method tailored for specific KBs. During the tuning stage, inspired by human learning strategies such as summarizatio and self-reflection, we propose a combined long- and short-dependency annotation method, as well as an iterative tuning approach. These techniques enable low-cost targeted training data augmentation and efficient adaptation without requiring external supervision. In the inference stage, we enhance the model’s performance on KBQA tasks using query expansion and sampling-based self-verify strategies. Our approach demonstrates significant improvements across various datasets spanning different domains and formats. Additionally, detailed analysis provides empirical guidance regarding the best data amount required.
@@ A.3 Supplementary Case
Figure 6: More cases for KBAda and baseline comparison.

### v3

# KBAlign: Efficient Self Adaptation on Specific Knowledge Bases

@@ Abstract
Humans can utilize techniques to quickly acquire knowledge from specific materials in advance, such as creating self-assessment questions, enabling us to achieving related tasks more efficiently. In contrast, large language models (LLMs) usually relies on retrieval-augmented generation to exploit knowledge materials in an instant manner, or requires external signals such as human preference data and stronger LLM annotations to conduct knowledge adaptation. To unleash the self-learning potential of LLMs, we propose KBAlign, an approach designed for efficient adaptation to downstream tasks involving knowledge bases. Our method utilizes iterative training with self-annotated data such as Q&A pairs and revision suggestions, enabling the model to grasp the knowledge content efficiently. Experimental results on multiple datasets demonstrate the effectiveness of our approach, significantly boosting model performance in downstream tasks that require specific knowledge at a low cost. Notably, our approach achieves over 90% of the performance improvement that can be obtained by using GPT-4-turbo annotation, while relying entirely on self-supervision. We release our experimental data, models, and process analyses to the community for further exploration  111 https://github.com/thunlp/KBAlign.
KBAlign: Efficient Self Adaptation on Specific Knowledge Bases
Zheni Zeng1*, Yuxuan Chen1*, Shi Yu1, Ruobing Wang3, Yukun Yan1🖂, Zhenghao Liu2, Shuo Wang1, Xu Han1, Zhiyuan Liu1, Maosong Sun1, 1Tsinghua University, Beijing, China 2Northeastern University, Shenyang, China 3University of Chinese Academy of Sciences, Beijing, China yanyk.thu@gmail.com
@@ 1 Introduction
Figure 1: Comparison between different knowledge adaptation methods. KBAlign requires no external signals for tuning, and improves by itself in both tuning and inference stages.
@@
Inspire by the above analogy, we purpose KBAlign, a highly efficient self adaptation approach tailored to specific knowledge bases (KBs) comprising self-improvement learning. Generally, we tune the model in a completely self-supervised manner, and maximize the tuned model’s potential in inference. As shown in Fig. 1, for self annotation, we organize the original KB materials and conduct self annotation to get initial Q&A data pairs that can cover various downstream task scenarios. For iterative tuning, we require the model to check its own responses as the learning progresses, and help the model to modify the common mistakes at the current stage. Meanwhile, we conduct targeted inference in which strategies such as query expansion and confidence verification are adopted to refine the response.
Experiments on fact QA, long-form QA, and professional field test have shown the effectiveness of our method across different backbone models. With a low cost, KBAlign masters the general knowledge content from the KB and achieves significant performance improvements. Side experiments including ablation studies and performance curve analyses identify the most efficient self-annotated data amount and optimal training volumes, offering valuable guidance for effectively applying KBAlign in practical scenarios.
Our main contributions are as follows: (1) We propose KBAlign, a novel method for autonomous LLM adaptation tailored to knowledge bases. KBAlign is the first approach to help LLMs perform KB adaptation relying entirely on self annotations; (2) We provide empirical insights into efficient self adaptation to KBs, offering practical parameters and settings for deploying KBAlign; (3) We conduct a comprehensive analysis of the proposed self adaptation framework. Through a range of evaluation metrics and case studies, we identify the effectiveness of KBAlign and discuss the current limitations of our approach, highlighting areas for future improvement.
@@ 2 Related Work
Figure 2: Details for the KBAlign framework. (a) Schematic and data composition for the whole process; (b) Instances for different annotation strategies and tasks.
@@ 4.3 Other Settings
Table 1: KBAlign adaptation experiments on LooGLE, ASQA, and JEC-QA. We report average for 3 random seeds.
@@ 4.4 Result Analysis
Main Experiments. The main results for our experiments are shown in Table 1. We provide the GPT-series results, the initial version and the self-adapted version of both MiniCPM-2.4B and LLaMA-3.1-8B-Instruction on the three dataset. Overall, comparing the “Ours” lines with the corresponding original model, we can see that KBAlign ensures an obvious improvement on most of the metrics, regardless of the dataset and the backbone model. This shows the generalization and effectiveness of our method.
@@
Figure 3: The impact of training amount on LooGLE performance. ‘w GPT‘ refers to training with GPT-annotated data.
@@
Figure 5: Cases for KBAlign and baseline comparison. We display the translation for the Chinese JEC-QA task. The bold text and underlined text providing correct and wrong information for the QA process.
@@ 4.5 Case Studies
We display typical cases in Fig. 5 to explain the specific usefulness of our strategies. “Baseline” refers to MiniCPM-2.4B and “ours” refers to the adapted version of it. Overall, KBAlign achieves a general grasp of current KB, a better knowledge answering, and a reasonable confidence verification.
@@ 5 Discussion and Limitations
In this paper, we introduced KBAlign, a highly efficient self-adaptive method tailored for specific KBs. During the tuning stage, inspired by human learning strategies such as summarizatio and self-reflection, we propose a combined long- and short-dependency annotation method, as well as an iterative tuning approach. These techniques enable low-cost targeted training data augmentation and efficient adaptation without requiring external supervision. In the inference stage, we enhance the model’s performance on KBQA tasks using query expansion and sampling-based self-verify strategies. Our approach demonstrates significant improvements across various datasets spanning different domains and formats. Additionally, detailed analysis provides empirical guidance regarding the best data amount required.
@@ A.3 Supplementary Case
Figure 6: More cases for KBAlign and baseline comparison.